What's the total mechanical energy of a 400kg satellite in circular orbit 1100km above Earth's surface?
Answer #1

助一下 G Mm
Answer #2


Total mechanical energy E = G*mE*m/2r

here r is radius of circular orbit =Radius of earth+1100 km = 6380+1100 = 7480 km
(6.67*10^-11*5.97*10^24*400)/(2*7.48*10^6) = 1.06*10^10 J
Answer #3

Use Kepler's third law: T